OpenStudy (anonymous):

evaluate each expression for the given values of the variables . 1. 4a + 6b + 7 for a = 2 and b = 3 2. 7y^2 + 7y for y = 3

OpenStudy (anonymous):

1. 4(2)+6(3)+7=8+18+7=33

OpenStudy (anonymous):

2. 7(3)^2 +7(3)=63+21=84